# Regularized Newton Raphson Inversion for Text-to-Image Diffusion Models

### v2

@@ 2 Related work
Alternative approaches proposed in [15, 4] suggest inverting with a DDPM scheduler instead of DDIM. They begin by constructing auxiliary images $x_{1},...,x_{T}$ and then extracting noise maps $z_{1},...,z_{T}$ to achieve error-free image reconstruction. Despite their effectiveness, the stochastic nature of these methods presents several challenges. Firstly, it necessitates storing $T+1$ latents for every inverted image, resulting in an additional $16K\times T$ parameters. Secondly, the method is solely applicable to reconstruction and editing tasks. Lastly, DDPM inversion typically yields lower-quality results compared to DDIM inversion  [14]. Since our focus is on DDIM deterministic inversion, these papers fall outside the scope of this paper.
@@ 3 Preliminaries
We first establish the fundamentals of Denoising Diffusion Implicit Models (DDIMs). In this model, a backward pass (denoising) is the process that generates an image from a seed noise. A forward pass is the process of adding noise gradually to an image until it becomes pure Gaussian noise. Inversion is similar to the forward pass but the goal is to end with a specific Gaussian noise that would generate the image if denoised.
Forward Pass in Diffusion Models. Latent diffusion and consistency models learn to generate images through a systematic process of iteratively adding Gaussian noise to a latent data sample until the data distribution is mostly noise. The data distribution is subsequently gradually restored through a reverse diffusion process initiated with a random sample (noise seed) from a Gaussian distribution. In more detail, the process of mapping a (latent) image to noise is a Markov chain that starts with $z_{0}$, and gradually adds noise to obtain latent variables $z_{1},z_{2},\dots,z_{T}$, following $q(z_{1},z_{2},\ldots,z_{T}|z_{0})=\Pi^{T}_{t=1}{q(z_{t}|z_{t-1}})$, where $\forall t:z_{t}\in\mathbb{R}^{d}$ with $d$ denoting the dimension of the space. Each step in this process is a Gaussian transition
$q(z_{t}|z_{t-1})\sim\mathcal{N}(z_{t},\sqrt{1-\beta_{t}}z_{t-1},\beta_{t}I),$ (1)
@@
Denoising Diffusion Implicit Models (DDIM). Sampling from diffusion models can be viewed alternatively as solving the corresponding diffusion Ordinary Differential Equations (ODEs) [18]. DDIM [32] scheduler, a popular deterministic scheduler, proposed denoising a latent noise vector in the following way:
@@
where $\psi(\alpha)=\sqrt{\frac{1}{\alpha}-1}$, and $\Delta\psi(\alpha_{t})=\psi(\alpha_{t})-\psi(\alpha_{t-1})$.
$\epsilon_{\theta}(z_{t},t,p)$ is the output of a network that was trained to predict the noise to be removed.
@@ 4 Our method: Regularized Newton Raphson Inversion
DDIM inversion above often yields poor reconstruction accuracy. Proposed improvements described above have a linear convergence rate and may take many seconds to compute. In this paper we describe a faster and more robust alternative based on the well-known Newton-Raphson method (NR) [16]. Newton-Raphson is a method for iteratively finding the roots of a system of equations.
@@
and seek for solutions that satisfy $\hat{r}(z_{t})=0$. Here, |||||$\cdot$||| | ⋅ | | denotes a norm; we used $L_{1}$ to simply sum over all absolute values of $\hat{r}$. This reduction transforms the Jacobian matrix into a vector, making it easy and fast for computation (see derivation in Appendix A). Under some conditions, and particularly when there is only a single root in the $\epsilon$-neighborhood of the initial guess, the Newton-Raphson method can be proved to have a quadratic convergence rate [16]. In fact, $\hat{r}$ introduces an underdetermined equation that may have multiple roots, and in practice, we find that solving Eq. 6 for $r(z_{t})=0$ often converges to a solution that is out-of-distribution for the diffusion model. This inversion solution leads to bad reconstruction. We address this issue in the next section.
@@ 4.1 Regularized Newton Raphson
$\mathcal{F}(z_{t}):=||z_{t}-f(z_{t})||_{1}-\lambda\sum_{j=1}^{d}\log q_{j}(z_{t}|z_{t-1})$ (7)
Here, $\lambda>0$ is a hyperparameter weighting factor for the regularization, and $q_{j}$ is the $j$’th component of $q(z_{t}|z_{t-1})$ (refer to Appendix D for more details). We now seek a solution that satisfies $\mathcal{F}(z_{t})=0$. Note that this equation is strictly satisfied iff the residual $\hat{r}(z_{t})=0$ and the probabilities $q_{j}(z_{t}|z_{t-1})=1$, for all $j$.

### v3

@@ 2 Related work
Alternative approaches proposed in [15, 4] suggest inverting with a DDPM scheduler instead of DDIM. They begin by constructing auxiliary images $x_{1},...,x_{T}$ and then extracting noise maps $z_{1},...,z_{T}$ to achieve error-free image reconstruction. Despite their effectiveness, the stochastic nature of these methods presents several shortcomings. Firstly, it necessitates storing $T+1$ latents for every inverted image, resulting in an additional $\sim 16K\times T$ parameters. Secondly, [10, 14] argue that the focus on noise correction to preserve the original image during the inversion process causes the edit-friendly DDPM to struggle with reconstructing fine details. Finally, DDPM requires approximately 100 steps for high-quality inversion, making it time-consuming, with each image edit taking around 30 seconds. In this paper, we focus on fast and accurate inversion methods, while comparing our approach to state-of-the-art deterministic inversion approaches.
@@ 3 Preliminaries
We first establish the fundamentals of Denoising Diffusion Implicit Models (DDIMs). In this model, a backward pass (denoising) is the process that generates an image from a seed noise. A forward pass is the process of adding noise gradually to an image until it becomes pure Gaussian noise. Inversion is similar to the forward pass but the goal is to end with a specific Gaussian noise that would generate the image if denoised. For more details see [32].
Forward Pass in Diffusion Models. Latent diffusion models [26] learn to generate images through a systematic process of iteratively adding Gaussian noise to a latent data sample until the data distribution is mostly noise. The data distribution is subsequently gradually restored through a reverse diffusion process initiated with a random sample (noise seed) from a Gaussian distribution. In more detail, the process of mapping a (latent) image to noise is a Markov chain that starts with $z_{0}$, and gradually adds noise to obtain latent variables $z_{1},z_{2},\dots,z_{T}$, following a distribution $q(z_{1},z_{2},\ldots,z_{T}|z_{0})=\Pi^{T}_{t=1}{q(z_{t}|z_{t-1}})$, where $\forall t:z_{t}\in\mathbb{R}^{d}$ with $d$ denoting the dimension of the space. Each step in this process is a Gaussian transition
$q(z_{t}|z_{t-1}):=\mathcal{N}(z_{t};\mu_{t}=\sqrt{1-\beta_{t}}z_{t-1},\Sigma_{t}=\beta_{t}I),$ (1)
@@
Denoising Diffusion Implicit Models (DDIM). Sampling from diffusion models can be viewed alternatively as solving the corresponding diffusion Ordinary Differential Equations (ODEs) [18]. DDIM [32], a popular deterministic scheduler, proposd to denoise a latent noise vector in the following way:
@@
where $\alpha_{t}=1-\beta_{t}$, $\psi(\alpha)=\sqrt{\frac{1}{\alpha}-1}$, and $\Delta\psi(\alpha_{t})=\psi(\alpha_{t})-\psi(\alpha_{t-1})$. $\epsilon_{\theta}(z_{t},t,p)$ is the output of a network that was trained to predict the noise to be removed.
@@ 4 Our method: Regularized Newton Raphson Inversion
The DDIM inversion method suggests a fast, yet inaccurate inversion (as observed from poor reconstructions). Previous Fixed Point iterations and Gradient Descent Methods suggest relatively fast inversions with a better accuracy than DDIM. Yet, these methods have a linear convergence rate that may take many iterations and long time to compute (many seconds). We introduce a faster and more accurate alternative, based on the well-known Newton-Raphson iterative method (NR) [16]. Newton-Raphson is a method for iteratively finding the roots of a system of equations.
@@
and seek for solutions that satisfy $\hat{r}(z_{t})=0$. Here, |||||$\cdot$||| | ⋅ | | denotes a norm; we used $L_{1}$ to simply sum over all absolute values of $\hat{r}$. This reduction transforms the Jacobian matrix into a vector, making it easy and fast for computation (see derivation in Appendix A). Under some conditions, and particularly when there is only a single root in the $\epsilon$-neighborhood of the initial guess, the Newton-Raphson method can be proved to have a quadratic convergence rate [16]. In fact, $\hat{r}$ introduces an underdetermined equation that may have multiple roots, and in practice, we find that solving Eq. 6 for $r(z_{t})=0$ often converges to a solution that is out-of-distribution in the diffusion model. This inversion solution leads to bad reconstruction. We address this issue in the next section.
@@ 4.1 Regularized Newton Raphson
$\mathcal{F}(z_{t}):=||z_{t}-f(z_{t})||_{1}-\lambda\log\tilde{q}(z_{t}|z_{t-1})$ (7)
Here, $\lambda>0$ is a hyperparameter weighting factor for the regularization (see appendix D for ablation on $\lambda$ values). While $q$ is a Gaussian distribution with a scalar covariance matrix (Eq.(1)), $\tilde{q}$, represents the exponent term of $q$. For a sampled $z_{t}$, $\log\tilde{q}(z_{t}|z_{t-1})$, can be computed efficiently by summing over its components. We now seek a solution that satisfies $\mathcal{F}(z_{t})=0$. Note that this equation is strictly satisfied iff the residual $\hat{r}(z_{t})=0$ and $z_{t}=\mu_{t}$ i.e. $z_{t}$ is at the center of the distribution.